OpenStudy (cutiecomittee123):

Use the trigonometric subtraction formula for sine to verify this identity sin(pi/2-x)=cosx First of all what is the trigonometric subrtraction formula? Second of all, how is this solved?

OpenStudy (superdavesuper):

subtraction formula for sine: sin(A-B)=sin A cos B - cos A sin B so A=pi/2 and B=x in this case...

OpenStudy (cutiecomittee123):

Okay so the left side then changes to sin pi/2 cos x - cos pi/2 sin x? So the whole equation is then sin pi/2 cos x - cos pi/2 sin x=cos x

OpenStudy (superdavesuper):

u got it...good job! :)

OpenStudy (cutiecomittee123):

but now what?

OpenStudy (princeharryyy):

you did a great job. you got the final answer.

OpenStudy (cutiecomittee123):

thats all? really

OpenStudy (princeharryyy):

subrtraction formula>> sin(A-B)=sin A cos B - cos A sin B sin(pi/2-x) = sin pi/2 cos x - cos pi/2 sin x as cos pi/2 = 0 and sin pi/2 = 1 expression becomes, sin(pi/2-x) = (1)*cos x - 0*sin x

OpenStudy (superdavesuper):

yes @cutiecomittee123 wat else u expecting? :P

OpenStudy (cutiecomittee123):

Well then if thats it can either of you help me with another one?

OpenStudy (princeharryyy):

and you have already done all that.

OpenStudy (cutiecomittee123):

or a few other ones lol

OpenStudy (princeharryyy):

how many r there?

OpenStudy (superdavesuper):

i will let @princeharryyy help u :) good luck @cutiecomittee123

OpenStudy (princeharryyy):

thanks man @superdavesuper

OpenStudy (princeharryyy):

u gave her the solution.

OpenStudy (cutiecomittee123):

Okay derive the trigonometric addition formula for sine sin(a+b)=sinacosb+cosasinb

OpenStudy (cutiecomittee123):

there are a few maybe 3 questions besides the one I just asked

OpenStudy (princeharryyy):

are you sure that you have to derive it, coz it's lengthy.